OpenStudy (anonymous):

What is the sum of the arithmetic sequence 3, 9, 15…, if there are 24 terms?

OpenStudy (anonymous):

+6

OpenStudy (anonymous):

1,728 1,452 2,028 2,268

OpenStudy (anonymous):

are my options

OpenStudy (anonymous):

1728 is the correct answer

OpenStudy (anonymous):

thanks soo much (:

OpenStudy (anonymous):

welcome. This is the approach I used: The series can be written as (6j-3) where j varies from 1 through 24 ( or any n) The summation of the series can therefore be written as ∑_(n=1)^n▒(6j-3) = 6 (∑j) – 3n =6*n(n+1)/2 -3n Substitute n with any number you need Substituting 24 for n yields the required result.